# Sufficiently Regularized Nonnegative Quartic Polynomials are Sum-of-Squares

### v1

@@ Abstract
Hilbert’s 17th problem famously established that not all nonnegative polynomials admit an SoS representation. Hilbert also identified a few special classes in which nonnegativity and SoS are equivalent, such as univariate polynomials, quadratic polynomials, and bivariate quartic polynomials. In this paper, we extend this equivalence to several new subclasses of multivariate quartically-regularized polynomials and characterise the NP-hardness boundary of these special-structure polynomials. Namely, we consider the global optimization of multivariate symmetric cubic polynomials regularised by weighted quartic powers of the Euclidean norm. These special-structure polynomials arise as iterative subproblems in high-order tensor methods for nonconvex optimisation problems. We consider shifting these polynomials by their global optimum so as to make them nonnegative, and show that for sufficiently high regularization and under mild assumptions, these polynomials admit a Sums-of-Squares (SoS) representation. We also identify several structured subclasses of quartically-regularized cubic polynomials for which global optimality of the model implies that its nonnegativity is certified by an SoS decomposition, for all values of the regularization parameter (such as quadratic-quartic polynomials and quartic polynomials containing a special cubic term that can be decomposed into a product of a quadratic norm and a linear form). We provide counterexamples using quartic separable norms that showcase the crucial role of the Euclidean norm in these representations. We illustrate how to use our SoS-based certificates for the Taylor subproblems inside high-order tensor methods for nonconvex optimization, with encouraging results.
1 Introduction
A foundational milestone in real algebraic geometry is Hilbert’s seminal 1888 paper [30], which provides a complete classification of the degrees and dimensions for which every nonnegative polynomial admits a representation as a sum of squares of polynomials. In this work [30], Hilbert showed that there exist nonnegative polynomials that cannot be expressed as sums of squares, although explicit examples of such polynomials were only constructed nearly eighty years later. The study of the gap between nonnegativity and sums of squares remains an active area of research to this day.
Motivated by new applications and growing interest in tensor methods and high-order optimization, there has been growing engagement within the optimization community in understanding when nonnegative polynomials can be represented as sums of squares (SoS). In particular, the minimization of regularized polynomials lies at the core of all regularization-based $p$th-order optimization methods ($p\geq 2$) and is fundamental to polynomial optimization. In these methods, we consider the unconstrained, potentially nonconvex, optimization problem $\min_{x\in\mathbb{R}^{n}}f(x)$, where $f:\mathbb{R}^{n}\to\mathbb{R}$ is $p$-times continuously differentiable and bounded below, for some $p\geq 2$. The adaptive $p$th-order regularization method (ARpp) is known to achieve optimal global and local convergence rates for sufficiently smooth objectives, in both convex and nonconvex settings [19, 16, 7, 49]. In the ARpp method, the subproblem is constructed using the $p$th-order Taylor expansion $T_{p,x_{k}}(s)$ of the objective at the iterate $x_{k}$, leading to the regularized polynomial,
$\min_{s\in\mathbb{R}^{n}}\;m_{p}(s):=T_{p,x_{k}}(s)+\frac{\sigma}{p+1}\|s\|^{p+1},$ (ARpp Subproblem)
where $\|\cdot\|$ denotes the Euclidean norm unless otherwise specified, while $\sigma>0$ is an adaptive regularization parameter. In [19, 9, 16, 15], it was proved that when $\sigma$ is sufficiently large – that is, when the model is adequately regularized – the step obtained from minimizing the subproblem is guaranteed to produce a decrease in the objective function value, a key ingredient for ensuring global convergence and rate of the ARpp method. In the nonconvex setting, Birgin et al [9] and Cartis et al [19, 16] have shown that the ARpp method requires at most $O\left(\epsilon_{g}^{-\frac{p+1}{p}}\right)$ evaluations of $f$ and its derivatives to obtain an iterate satisfying $\|\nabla f(x_{k})\|\leq\epsilon_{g}$, and that this bound is optimal within the class of smooth objectives with Lipschitz continuous $p$th order derivatives [15, 16]. These bounds further improve when $f$ is convex, as shown by Nesterov [49] and others. Moreover, ARpp methods achieve $p$th order local rates of convergence faster local convergence rates than first- and second-order methods [24, 71]. This demonstrates that increasing the model degree and accuracy order to $p>2$ leads to improved evaluation complexity and local rates compared to first- and second-order methods, providing strong motivation for the interest and use of higher-order methods in practice [20]. However, the question of how to efficiently solve the subproblems in ARpp methods in ways that exploit problem structure has only mainly been addressed in the convex case (see Nesterov [49, 53, 2] and follow-up works) until recently [74, 21] and very much remains an ongoing topic of research, especially in terms of understanding the level of computational difficulty involved.
It is well known that for polynomials of degree four or higher, deciding nonnegativity is NP-hard (see, for example, the NP-hardness of testing matrix copositivity [44]). Also, Nesterov has shown that maximizing a cubic polynomial over the unit sphere is also NP-hard [12, 39]. Although these results illustrate the general intractability of nonnegativity for high-degree polynomials, regularized polynomials form a more subtle and structured subclass. Interestingly, Parrilo et al. [60, Sec. 5.1] observed numerically that certain quartically regularized test functions appear to become SoS once the regularization parameter is chosen large enough. However, these observations were empirical; no theoretical proof, nor any explicit bound on the regularization magnitude, have been established in the literature. Recently, Ahmadi et al. [1] proved that a sufficiently regularized locally convex polynomial is SoS-convex, and the algorithmic framework is given in [75]. However, when (ARpp Subproblem) is not positive semidefinite, the presence or absence of the SoS property for the ensuing regularised polynomial model is an open question in the literature. The question of whether a regularized (nonnegative) even-degree symmetric polynomial necessarily becomes SoS for large regularization values remains essentially unexplored.
In the case of $p=3$, we let $f_{0}:=f(x_{k})\in\mathbb{R}$, $g:=\nabla f(x_{k})\in\mathbb{R}^{n}$, $H:=\nabla^{2}f(x_{k})\in\mathbb{R}^{n\times n}$ , $T:=\nabla^{3}f(x_{k})\in\mathbb{R}^{n\times n\times n}$. Then the (ARpp Subproblem) becomes a quartically-regularized cubic polynomial in $n$ variables,
$\min_{s\in\mathbb{R}^{n}}m_{3}(s)=f_{0}+g^{T}s+\frac{1}{2}H[s]^{2}+\frac{1}{6}\,T[s]^{3}+\frac{\sigma}{4}\,\|s\|^{4},$ (Quartically Regularized Cubic Model)
where $f_{0}$, $g$, $H$, and $T$ denote the constant, linear, quadratic, and cubic coefficients, respectively. Throughout, we assume that $H$ and $T$ are symmetric. Note that $H[s]^{2}$ (equivalently, $s^{T}Hs$) and $T[s]^{3}$ represent the second- and third-order directional derivatives at $x_{k}$ along $s\in\mathbb{R}^{n}$, respectively. The regularization term is a quartic-degree Euclidean norm, that is, $\frac{\sigma}{4}\|s\|^{4}=\frac{\sigma}{4}\,(s_{1}^{2}+\dotsb+s_{n}^{2})^{2}.$ Since $\sigma>0$, the model $m_{3}(s)$ is bounded below and the optimization problem is well defined. We denote the global minimum value of $m_{3}$ by $m_{*}$, and the global minimum is attained (see Remark 2.3). In this work, we aim to address the following fundamental question
When (Quartically Regularized Cubic Model) is sufficiently regularized (i.e., when $\sigma>0$ is large enough), is the shifted polynomial $m_{3}-m_{*}$ a sum of squares (SoS)?
Once a polynomial can be written as a sum of squares, this not only certifies its global nonnegativity but also allows the associated global polynomial optimization problem to be reformulated as a semidefinite program (SDP) at the zeroth level of the Lasserre hierarchy [33, 60]. Such SDPs have polynomial size $\binom{n+4}{4}$ and can be solved in polynomial time using, for example, interior-point methods [45]. Moreover, if the global minimizer is unique, it can be extracted in polynomial time through the moment–SoS hierarchy [36, 34].
In this work, we characterize the NP-hardness boundary of quartically regularized polynomials, thereby extending Hilbert’s classical characterizations to a new subfamily of multivariate polynomials. The contributions of the paper are as follows.
Regularization threshold for SoS: Under mild assumption, we establish an explicit bound on the regularization parameter $\sigma$ such that, if $\sigma$ exceeds this threshold, the shifted quartically regularized polynomial $(m_{3}-m_{*})$ is SoS. To the best of our knowledge, this is the first work to establish SoS results for a generally nonconvex quartically-regularized cubic polynomial; thus, building an understanding of how strength of regularization governs convexification, SoS certificates, and global optimality.
Special polynomial classes with equivalence between global optimality and SoS certificates: Hilbert’s 17th problem famously established that not all nonnegative polynomials admit an SoS representation [30]. Hilbert also identified a few special classes in which nonnegativity and SoS are equivalent, such as univariate polynomials, quadratic polynomials, and bivariate quartic polynomials. In this paper, we extend this equivalence to several new variants of multivariate quartically regularized polynomials. Specifically, (i) quadratic–quartic polynomials with any number of variables, and (ii) quartic polynomials containing a special cubic term that can be decomposed into a product of a quadratic norm and a linear form. These cases constitute a rare subclass of nonconvex polynomials that are SoS-certifiable, establishing the relationship between global optimality conditions and SoS representability.
Discussion of quartic norms: Lastly, we identify a subtle yet important distinction between different types of regularization norms. Several alternative regularization norms and seminorms have been proposed in the literature. For instance, Schnabel et al. [66] introduced a quartic seminorm formulation for unconstrained optimization of the form $\frac{1}{4}\sum_{j=1}^{k}\sigma_{j}(a_{j}^{T}s)^{4}$, where $\sigma_{j}>0$ and $a_{j}\in\mathbb{R}^{n}$. [76] discusses the separable quartic norm $\|s\|_{4}^{4}=\sum_{i=1}^{n}s_{i}^{4}$. We use explicit examples to illustrate that when the regularization term is not the Euclidean quartic norm $\|s\|^{4}$ but the separable quartic norm $\|s\|_{4}^{4}$, even for arbitrarily large $\sigma$, the corresponding nonnegative separable quartic polynomial need not be SoS. We provide detailed characterization and examples on the choice of regularization norm that leads to SoS representability.
The paper is organized as follows. Section 1.1 provides background and definitions on the SoS framework, the Lasserre hierarchy, and related results by Ahmadi, Lasserre, Nesterov, and Parrilo. Section 2 presents our main theoretical results, including explicit bounds on $\sigma$ under which the quartically regularized cubic polynomial becomes SoS. We further discuss how this bound differs between locally convex cases ($H\succ 0$) and locally nonconvex cases. Section 3 analyzes several special classes of quartically regularized polynomials, identifying which forms are SoS and which are not, and highlighting the distinctions arising from different regularization norms. Section 4 provides detailed examples and preliminary numerical illustrations illustrating how various regularization norms lead to different NP-hardness behaviors. Finally, Section 5 concludes the paper.
1.1 Preliminaries and Background on Nonnegative and SoS Polynomials
@@
Let $\mathcal{P}_{n}^{\deg(h)}$ denote the class of real polynomials $h(s):\mathbb{R}^{n}\to\mathbb{R}$ with $n$ variables and degree $\deg(h)$ and let $h(s):\mathbb{R}^{n}\to\mathbb{R}$ be a (multivariate) real polynomial in variables $s=[s_{1},\dotsc,s_{n}]^{T}\in\mathbb{R}^{n}$. The degree of a polynomial, $\deg(h)$, is defined to be the highest degree of its component monomials [^1] The degree of a monomial $s^{\alpha}$ where $\alpha=[\alpha_{1},\dotsc,\alpha^{n}]$ is equal to $\alpha_{1}+\dotsc+\alpha_{n}$. If the global minimum exists and is attained, we denote it by $h^{*}:=\min_{s\in\mathbb{R}^{n}}h(s)$ and $s^{*}$ is a global minimizer. A homogeneous polynomial is a polynomial where all the monomials have the same degree. A polynomial $h(s)$ is said to be nonnegative if $h(s)\geq 0$ for all $s\in\mathbb{R}$.
Definition 1.1.
@@
(SoS polynomial [1, 5, 32]) $h(s)$ is SoS if there exist polynomials $\tilde{h}_{1},\dotsc,\tilde{h}_{r}:\mathbb{R}^{n}\rightarrow\mathbb{R}$, for some $r\in\mathbb{N}$, such that $h(s)=\sum_{j=1}^{r}\tilde{h}_{j}(s)^{2}$ for all $s\in\mathbb{R}^{n}$. We refer to a problem as being SoS-exact at level 0 if $h(s)-h^{*}$ is a sum of squares polynomial.
Definition 1.1 implies that the degree of an SoS polynomial $h(s)$ has to be even, $\deg(h)=2d$, and that the maximum degree of each $\tilde{h}_{j}$ is $\frac{\deg(h)}{2}=d$. Let $\mathbb{R}[s]^{n\times n}$ be the real vector space of $n\times n$ real polynomial matrices, where each entry of such a matrix is a polynomial with real coefficients.
Definition 1.2.
(SoS-convex, [5, Def. 2.4]) $h(s)$ is SoS-convex if its Hessian $\nabla^{2}h(s)$ is an SoS-matrix. A symmetric polynomial matrix $\check{H}(s)\in\mathbb{R}[s]^{n\times n}$ is an SoS-matrix if there exists a polynomial matrix $\check{V}(s)\in\mathbb{R}[s]^{n_{1}\times n}$ for some $n_{1}\in\mathbb{N}$, such that $\check{H}(s)=\check{V}(s)^{T}\check{V}(s)$.
Clearly, if $h(s):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is SoS-convex, then $h(s)$ is convex [5, 32]. Recently, Ahmadi et al. [1] proved that for a locally strongly convex quartically regularized polynomial ( $H\succeq\delta I_{n}\succ 0$), there exists a sufficiently large regularization parameter $\sigma$ such that the polynomial becomes SoS-convex (see Theorem 1.1).
Theorem 1.1.
[…]
Lemma 1.1.
While convexity often leads to tractable global optimization [59, 5, 49], local nonconvexity makes the problem markedly more difficult. In 1992, the question of determining whether a polynomial defines a convex function was posed as one of the seven major open problems in the complexity theory of numerical optimization [58]. Ahmadi, Olshevsky, and Tsitsiklis later proved that this decision problem is strongly NP-hard even for quartic polynomials [4]. A natural and fundamental question, first posed and (partially) answered by Hilbert, is to determine for which dimensions $n$ and degrees $\deg(h)$, nonnegative polynomials or forms can always be expressed as sums of squares.
Theorem 1.2 (Hilbert, [30]).
For the space of real polynomials $\mathcal{P}_{n}^{\deg(h)}$, nonnegativity and sum-of-squares coincide if and only if one of the following holds:
  1. 1.
    $n=1$ (univariate polynomials of any degree);
  2. 2.
    $\deg(h)=2$ (quadratic polynomials in any number of variables);
  3. 3.
    $n=2$ and $\deg(h)=4$ (bivariate quartic polynomials).
For all other combinations of dimension and degree, Hilbert showed that nonnegativity does not imply a sum-of-squares representation. However, within the class of quartic polynomials, the SoS behavior of quartically regularized models is largely unexplored. Explicit examples of nonnegative but non-SoS polynomials appeared in the 1960s, most notably the Motzkin polynomial $M(s_{1},s_{2})=s_{1}^{4}s_{2}^{2}+s_{1}^{2}s_{2}^{4}-3s_{1}^{2}s_{2}^{2}s_{3}^{2}+s_{3}^{6},$ a nonnegative trivariate polynomial of degree six [43]. A more recent example appears in  [2, Thm. 3.3], a nonnegative trivariate quartic polynomial which is not SoS,
$\tilde{m}_{\mathcal{A}}(s)=\frac{9}{4}+2\|s\|^{2}+8(s_{1}s_{2}+s_{1}s_{3}+s_{2}s_{3})+(s_{1}^{4}+s_{2}^{4}+s_{3}^{4}).$ (1)
Using this example and related constructions, Ahmadi [2] proved that testing nonnegativity of separable quartic–quadratic polynomials is NP-hard. Note the distinction between (1) and (Quartically Regularized Cubic Model): the former uses a separable quartic norm regularization, whereas the latter employs a Euclidean quartic norm. This leads to subtle but important differences in the characterization of their SoS representations. Also, Nesterov [46] proved that even very structured nonconvex polynomial optimization problems are NP-hard to solve globally. In particular, he showed that the following problems are NP-hard.
Theorem 1.3.
([46, Thm 4], [39]) The maximization of homogeneous cubic or quartic form over the Euclidean ball, and the optimization of a quartic homogeneous polynomial under quadratic constraints are NP-hard [^2] The problem formulations are given in (52)–(54)..
A common feature among these NP-hard optimization problems in Footnote 2 is that they use constraints—such as Euclidean ball constraints, trust-region constraints, or quadratic constraints. In all three cases, intractability arises when the global minimizer is on the boundary of the feasible set, i.e., the constraint is active. In contrast, the regularization framework in (ARpp Subproblem) adopts a different globalization strategy by replacing explicit constraints with a regularization term. This guarantees boundedness from below and yields a well-defined unconstrained optimization problem. We show in later sections that, when the regularization parameter is sufficiently large, minimizing (Quartically Regularized Cubic Model) becomes tractable.
1.1.1 Connection to SDP and Matrix Generalizations
As mentioned earlier, the computational advantage of sums of squares is the connection to semidefinite programming. It is well known that a polynomial is a sum of squares if and only if it admits a positive semidefinite Gram matrix representation.
@@ 1.1.1 Connection to SDP and Matrix Generalizations
Example 1.1.

### v2

@@ Abstract
A polynomial that is nonnegative need not be a sum of squares of polynomials. This classical gap, identified by Hilbert in 1888, lies at the heart of why the global optimization of multivariate quartic polynomials is NP-hard. Yet we show that this gap is closed when using (sufficient) regularization, which fundamentally alters the algebraic structure of the problem. Namely, we investigate a class of quartically-regularized cubic polynomials which arise naturally in polynomial optimization and higher-order tensor methods for nonconvex problems. We show that, under mild assumptions and for sufficiently large Euclidean quartic regularization, the shifted nonnegative polynomial becomes a sum of squares, yielding an exact semidefinite programming (SDP) formulation at the zeroth level of the Lasserre hierarchy. We further derive explicit bounds on the regularization parameter that guarantee this property. Beyond this asymptotic regime, we identify structured subclasses for which SoS exactness holds for all regularization levels, including quadratic–quartic models and a class of low-rank-type cubic tensors. In contrast, we show that separable quartic regularized polynomials—including classical tensor models proposed by Schnabel (1991)—do not, in general, induce SoS representations, even under arbitrarily large regularization. Our results reveal a sharp structural boundary between tractable and intractable regimes in polynomial optimization. In particular, they explain why Euclidean quartic regularization plays a significant role: in addition to regularising the model, it can induce exact SoS certificates and exact SDP representations.
1 Introduction and Overview
A foundational result in real algebraic geometry, due to Hilbert (1888), shows that nonnegativity and sum-of-squares (SoS) representations do not coincide in general [22]. Outside a few special cases, there exist nonnegative polynomials that cannot be written as sums of squares. This gap lies at the heart of polynomial optimization and semidefinite programming.
From an optimization perspective, this distinction is fundamental. If a nonnegative polynomial admits an exact SoS certificate, then its global minimum can be computed via a single semidefinite program (SDP) at the zeroth level of the Lasserre hierarchy [24, 40]. By contrast, in the absence of such a certificate, one typically requires a hierarchy of increasingly large SDP relaxations. Although finite convergence can occur under suitable conditions [39], the required relaxation order is not known a priori and may be arbitrarily large, and the problem remains NP-hard in general. Classical examples illustrate this gap clearly. Nonnegative polynomials such as the Motzkin and Choi–Lam forms are not SoS [30, 18], and more recently, Ahmadi [2, Thm. 3.3] constructed an explicit nonnegative trivariate quartic polynomial that is not SoS. Furthermore, Nesterov [31] proved NP-hardness results for several cubic and quartic polynomial optimization problems over the Euclidean unit ball.
At the same time, many polynomial models arising in optimization are not arbitrary. In particular, structured regularized polynomials appear naturally in higher-order optimization methods and tensor-based models. These methods construct local Taylor approximations of an objective function and coupled with higher-degree regularization terms, leading to polynomial subproblems of increasing order (often referred to as ARpp models). In this paper, we focus on quartically regularized cubic polynomials of the form
$m_{3}(s)=f_{0}+g^{T}s+\frac{1}{2}H[s]^{2}+\frac{1}{6}\,T[s]^{3}+\frac{\sigma}{4}\,\|s\|^{4},$ (Quartically Regularized Cubic Model)
where $f_{0}\in\mathbb{R}$, $g\in\mathbb{R}^{n}$, $H\in\mathbb{R}^{n\times n}$, and $T\in\mathbb{R}^{n\times n\times n}$ are the constant, linear, quadratic, and cubic coefficients, respectively, and $\sigma>0$ is a regularization parameter; the quartic term uses the Euclidean norm $\|s\|^{4}=(s_{1}^{2}+\cdots+s_{n}^{2})^{2}$. When the minimization of $m_{3}$ appears as an iterative subproblem in optimization algorithms for minimizing a general objective, the parameter $\sigma$ is typically chosen adaptively to control the quality of the local model and its minimizers. When $\sigma$ is sufficiently large, the model minimization is guaranteed to yield a descent step for a general objective, which is a key ingredient for ensuring global convergence of adaptive regularization methods [14, 6, 11, 10]. Not only does (Quartically Regularized Cubic Model) appear naturally in polynomial optimization and higher-order methods, but can also be studied independently as a structured class of nonconvex polynomials.
For any $\sigma>0$, the model is bounded below and admits a well-defined global minimum value $m_{*}$, so that the quartic polynomial $m_{3}(s)-m_{*}$ is nonnegative for $s\in\mathbb{R}^{n}$. The central question of this paper is the following:
Under what conditions on the regularization weight $\sigma$ and problem data ($g$, $H$ and $T$) does the shifted polynomial $m_{3}-m_{*}$ admit an exact sum-of-squares representation?
At first glance, this may appear unlikely given the NP-hardness of general quartic polynomial optimization. However, empirical observations and recent theoretical developments suggest that regularization may induce hidden structure. In particular, Parrilo [40, Sec. 5.1] observed numerically that certain quartically regularized models appear to become SoS for sufficiently large regularization. More recently, Ahmadi et al. [1] proved that sufficiently regularized locally strongly convex quartic polynomials become SoS-convex, a related notion to SoS. However, in the general case – namely, when (Quartically Regularized Cubic Model) is nonconvex, whether $m_{3}-m_{*}$ admits an exact SoS representation for sufficiently large $\sigma$ remains an open question. Similarly open is the question of whether specific values of $g$, $H$ and $T$ yield SoS-representable polynomials for any value of $\sigma>0$. We address both of these questions here.
The key message of our paper is that the Euclidean quartic regularization in the definition of $m_{3}$ creates a structured, tractable regime, when chosen appropriately. While general nonnegative quartic polynomials need not be SoS, we show that sufficiently large regularization in (Quartically Regularized Cubic Model) induces sufficient structure in the polynomial so that an exact SoS certificate can be constructed. At the same time, this behavior is not universal across quartic regularizations: certain separable quartic regularization can fail to produce SoS exactness even for arbitrarily large regularization.
The main contributions of the paper are as follows.
(i) SoS exactness under sufficient regularization. We show that, under mild assumptions, choosing sufficiently large Euclidean quartic regularization in (Quartically Regularized Cubic Model) guarantees that the shifted polynomial $m_{3}-m_{*}$ is SoS. We derive explicit bounds on the regularization parameter $\sigma$ ensuring this property. To the best of our knowledge, this is the first SoS exactness result for generally nonconvex quartically regularized cubic polynomials. This shows that appropriate regularization can fundamentally alter SoS certifiability and enable exact semidefinite (SDP) representations.
(ii) Structured polynomial classes with exact SoS certificates. Beyond the large-regularization regime, we identify nontrivial subclasses for which SoS exactness holds for all $\sigma>0$. These include quadratic–quartic models and structured cubic tensors of the form $(t^{T}s)\|s\|^{2}$ (for given $t\in\mathbb{R}^{n}$ and for all $s\in\mathbb{R}^{n}$). These results extend Hilbert-type equivalences between nonnegativity and SoS to new families of multivariate polynomials.
(iii) A structural boundary: Euclidean vs. separable regularization. We show that the Euclidean quartic norm plays a critical role in enabling SoS exactness. In contrast, separable quartic regularization—including classical tensor models proposed by Schnabel [43]—may fail to produce SoS representations even for arbitrarily large regularization. This reveals a sharp boundary between tractable and intractable regimes in polynomial optimization. These findings also provide insight into higher-order optimization and tensor methods: Euclidean quartic regularization not only regularizes the subproblem, but also induces a regime where exact SoS certificates and thus exact SDP formulation is attainable.
The remainder of the paper is organized as follows. Section 1.1 reviews the necessary background on SoS polynomials, SoS-convexity, the Lasserre hierarchy, and related work in polynomial optimization. Section 2 contains the main theorem, proving that sufficiently strong Euclidean quartic regularization forces SoS exactness for the quartically regularized cubic model. Section 3 studies several structured subclasses, including cases where SoS exactness holds for all $\sigma>0$, and contrasts them with separable quartic regularization models where SoS exactness may fail. Section 4 presents examples illustrating why the geometry of the regularization term is decisive, and explains how our results differ from classical NP-hardness examples. Finally, Section 5 concludes with a discussion of the implications for higher-order optimization, tensor methods, and future directions.
1.1 Background: Nonnegativity, Sum of Squares, and SDP
@@
Let $\mathcal{P}_{n}^{\deg(h)}$ denote the class of real polynomials $h:\mathbb{R}^{n}\to\mathbb{R}$ of degree $\deg(h)$. We denote by $h^{*}:=\min_{s\in\mathbb{R}^{n}}h(s)$ the global minimum when it exists. A polynomial $h(s)$ is said to be nonnegative if $h(s)\geq 0$ for all $s\in\mathbb{R}^{n}$.
Definition 1.1 (SoS polynomial).
@@
A polynomial $h(s)$ is a sum of squares (SoS) if there exist polynomials $h_{1},\dots,h_{r}$ such that $h(s)=\sum_{j=1}^{r}h_{j}(s)^{2}.$ We say that $h$ is SoS-exact at level 0 if $h(s)-h^{*}$ is SoS.
Hilbert [22] gives that nonnegativity and SoS coincide only in limited cases (e.g., univariate polynomials or quadratic forms), while in general they differ. A key advantage of SoS representations is their connection to semidefinite programming (SDP). It is well known that a polynomial is a sum of squares if and only if it admits a positive semidefinite Gram matrix representation. This allows global minimization of SoS polynomials via a single SDP.
@@
If $h(s)-h^{*}$ is SoS, then its global minimum can be computed by solving a single semidefinite program via its Gram matrix representation [25, 40]. Moreover, under mild conditions (e.g., uniqueness), the global minimizer can be recovered from the solution of the corresponding moment SDP [38]. When $h(s)-h^{*}$ is not SoS, one can instead use the Lasserre hierarchy of SDP relaxations, which converges asymptotically and, under suitable conditions, in finitely many steps [24, 39]. However, the required relaxation order is not known a priori and may be arbitrarily large, so the problem remains NP-hard in general [42]. We will also use the notion of SoS-convexity. A polynomial $h$ is called SoS-convex if its Hessian is an SoS matrix. In this case, nonnegativity together with first-order optimality implies an SoS representation [20]. Recently, Ahmadi et al. [1] showed that sufficiently regularized locally strongly convex quartic polynomials become SoS-convex [^1] More details in SoS convex are given in Section B.1., with an algorithmic framework developed in [52], highlighting the tractability of the convex setting. Helton and Nie [20] showed that convex semialgebraic sets satisfying suitable curvature or SoS-type conditions admit SDP representations. However, when the model is nonconvex, the SoS behavior of quartically regularized polynomials remains largely unexplored.
Related work on high-order Methods
There is a substantial body of work on higher-order optimization methods, in both convex and nonconvex settings. The adaptive $p$th order regularization framework (ARpp) has been extensively studied in [14, 6, 11, 10]. In particular, it is known that sufficiently large regularization ensures descent and guarantees optimal evaluation complexity. In the nonconvex setting, ARpp methods require at most $O\left(\epsilon_{g}^{-\frac{p+1}{p}}\right)$ evaluations to achieve an $\epsilon_{g}$ approximate first order critical point and is optimal for this class of problems [6, 14, 11, 10]. Improved complexity bounds are available in the convex case [33], and higher-order methods also possess faster local convergence rates [19, 48]. These results highlight the practical and theoretical advantages of using higher-order models [15]. On the iterative algorithmic side, Nesterov and collaborators developed efficient solvers for convex higher-order models based on Bregman techniques and higher-order proximal methods [33, 34, 35, 36, 37, 32]. In the nonconvex setting, recent approaches exploit structured quartic regularization, including quadratic–quartic regularization (QQR) [51], cubic–quartic regularization (CQR) [16], and the diagonally tensor method (DTM) [53]. These methods leverage higher-order structure to design tractable subproblem solvers.
From a polynomial optimization perspective, Ahmadi and Zhang [5] established polynomial-time checkable conditions for local minimality of cubic polynomials. Building on this, Silina et al. [45] proposed an unregularized third-order Newton method that computes local minimizers of $m_{3}$ with $\sigma=0$, using SDPs, with a follow-up global algorithmic framework in [8]. A broader overview of higher-order methods can be found in [11, Ch. 4].
2 Regularization-Induced Sum-of-Squares Exactness
While general nonnegative quartic polynomials need not admit sum-of-squares (SoS) representations, we show that sufficiently large quartic regularization fundamentally alters this picture, inducing sufficient change in structure to make $m_{3}-m_{*}$ SoS representable.
We now explain the methodology underlying our main result. We first derive a constructive SoS representation based on global optimality conditions for (Quartically Regularized Cubic Model). We then quantify how the size of the minimizer scales with the regularization parameter. Finally, we show that the regularization term eventually dominates all remaining terms, ensuring that the resulting representation of $m_{3}-m_{*}$ is indeed SoS.
The two central results of this section are Theorem 2.1 and Theorem 2.3. The first gives a constructive SoS certificate conditional on $B(s^{*})\succeq 0$, while the second shows that this condition is automatically satisfied once $\sigma$ is sufficiently large. We also discuss two important regimes separately. In the locally strongly convex case, sufficiently large regularization leads to global convexification and SoS-convexity; see Section 2.3.1. In the locally nonconvex case, the model remains nonconvex near the origin, but we show that sufficiently large regularization still drives all relevant stationary points into a region where global optimality and SoS exactness can be certified; see Section 2.3.2. These two regimes together clarify how regularization, convexification, SoS certificates, and global optimality interact in quartically regularized cubic models.
2.1 A Constructive SoS Representation via Global Optimality Conditions
We begin by deriving a constructive sufficient condition under which the shifted polynomial $q(v):=m_{3}(s^{*}+v)-m_{3}(s^{*})$ is a sum of squares. The strategy is to rewrite $q(v)$ so that its quartic and cubic terms can be reorganized into explicit square expressions plus one residual quadratic form in $v$. The latter will be absorbed by a symmetric matrix $B(s^{*})$. If this matrix is positive semidefinite, then the whole polynomial becomes SoS.
The first step is an exact expansion of the shifted model around $s^{*}$. Using the global optimality conditions of $m_{3}$ in [53], we obtain the following representation. An equivalent integral representation of the shifted model and its connection to global optimality conditions are provided in Appendix C.1.
Lemma 2.1.
(Adapted from [53, Lemma 2.1]) Let $q(v):=m_{3}(s^{*}+v)-m_{3}(s^{*})$ where $m_{3}$ is defined in (Quartically Regularized Cubic Model). The first- and second-order derivatives of $m_{3}$ are
where $I_{n}\in\mathbb{R}^{n\times n}$ is the identity matrix. Then the fourth-order Taylor expansion of $q(v)$ at the point $s^{*}$ is exact and yields, for any vector $v\in\mathbb{R}^{n}$,
$q(v)=\nabla m_{3}(s^{*})^{T}v+\frac{1}{2}\,\nabla^{2}m_{3}(s^{*})[v]^{2}+\frac{1}{6}\,T[v]^{3}+\frac{\sigma}{4}\|v\|^{4}+\sigma\,{s^{*}}^{T}v\,\|v\|^{2}.$ (1)
Definition 2.1.